what is the mean for the combined set if the first sample has n=3 and the second sample has n=6
1
Expert's answer
2019-09-16T10:40:40-0400

"M=\frac{M_1n_1+M_2n_2}{n_1+n_2}=\frac{6*3+12*6}{3+6}=10"